OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. 25 4 0 1

OpenStudy (ash2326):

@SavannahChey can you find f(g(x))?

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

Work out g(2) and put it for x in f(x)

OpenStudy (ash2326):

We have \[f(x)=3x+1\] and g(x)=2x-4 so \[f(g(x))=3g(x)+1\] plugin g(x)=2x-4 to find f((g(x))

OpenStudy (anonymous):

Break it apart: g(3)= 3(3)-4=9-4=5. Thus, f(g(3))=f(5)=2(5)+1=10+1=11

OpenStudy (anonymous):

It's g(2) (zero)

OpenStudy (anonymous):

oh okay muh bad... ill fix it

OpenStudy (anonymous):

@SavannahChey got it?

OpenStudy (anonymous):

i think .. ;D so my answer is 1 , right ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks guys : )

OpenStudy (anonymous):

sorry for the confusion on my post... i thought it was g(3) not g(2)

OpenStudy (anonymous):

its okay .

OpenStudy (anonymous):

kk